OpenStudy (anonymous):

solve for x: x^2-3x-10=0

OpenStudy (luigi0210):

Factor it out: (x-5)(x+2) Set it equal to zero: (x-5)(x+2)=0 Solve: x-5=0 and x+2=0 So x= -2, and 5

OpenStudy (anonymous):

you r good at this